# Survey in Characterization of Semantic Change

### v4

@@ Abstract
Living languages continuously evolve to reflect the cultural changes of human societies. This evolution manifests through neologisms (new words) or the semantic change of existing words (new meanings for existing words). Understanding the meaning of words is vital for interpreting texts from different cultures (regionalisms or slang), domains (e.g., technical terms), or time periods. In computer science, this phenomenon is relevant to computational linguistics tasks such as machine translation, information retrieval, and question answering. Semantic change can impact the performance of these applications, making it important to understand and characterize these changes formally. This problem has recently attracted significant attention from the computational linguistics community. Several approaches can detect semantic changes with good precision, but more effort is needed to characterize how word meanings change and to determine how to mitigate the impact of this phenomenon. This survey provides a comprehensive overview of existing approaches to the *characterization of semantic change*. We also formally define three classes of characterization: change in dimension (whether a word’s meaning becomes broader or narrower), change in orientation (whether a word acquires a more pejorative or ameliorative sense), and change in relation (whether a word is used in a new figurative context, such as a metaphor or metonymy). We demonstrate the applicability of this formalism on existing corpora, summarize the key aspects of selected publications, and discuss current needs and trends in research on semantic change characterization.
@@ 1 Introduction
Language is one of the most dynamic skills humans have acquired, constantly evolving to adapt communication to new purposes and needs (pinker2003language, p. 13-23). While no single unified theory explains its origins (allan2013oxford, p. 13-52), language has clearly undergone significant modifications throughout human history. One of the prevailing theories of language evolution is based on the assumption of “continuity of change” (rilling2012continuity).
This theory posits that language, being too complex to have appeared suddenly, must have evolved gradually from earlier pre-linguistic systems used by our ancestors (Pinker1990NaturalLA, p. 725-726). In this context, comprehending language change contributes to the fields of history, sociology, linguistics, philosophy, and Natural Language Processing (NLP) (campbell2013historical, p. 1-8), as it is essential to understanding the origins of language, cultural development, grammatical theories, and the relationship between language and thought (Sherwood2008ANH, p. 426–454).
@@
  • Phonological Changes: Changes in the pronunciation and sound systems of a language. For instance, the pronunciation of vowel sounds changed significantly due to the Great Vowel Shift in Medieval English (nevalainen2016oxford, p. 367).
@@
  • Syntactic Changes: Changes in grammar and word order. For example, Modern English emphasizes word order and auxiliary verbs more than the intricate inflectional system of Old English (nevalainen2016oxford, p. 616).
@@
While phonological and morphological changes are crucial, semantics plays a pivotal role in language evolution. Semantic change reflects how words adapt to evolving cultural, technological, and societal contexts, creating new meanings and preserving language’s fundamental purpose of effective communication. Despite being a recurring topic in scientific research, semantic change remains less understood than other forms of linguistic evolution due to its often subtle nature compared to more observable shifts, like those resulting from grammar reforms (allan2012current, p. 2-3).
@@
The advent of computational linguistics and advanced models marked a major shift in the field. With vast digital corpora and efficient algorithms, researchers began to apply statistical methods from natural language processing to analyze text automatically. This marked a transition from purely observational methods to data-driven statistical analysis (allan2012current, p. 161-164).
With access to large text corpora and powerful algorithms, linguists can now detect subtle patterns in language with greater accuracy and efficiency. This allows for the study of languages with limited prior linguistic knowledge (Conneau2017WordTW; Wu2019BetoBB), given sufficient data, and enables comparison of multiple languages to test for universal properties in latent space (Vulic2020AreAG; Wang2022EnglishCL).
By adopting a model-based approach, linguists can now explore complex language phenomena across different cultures and backgrounds (Nivre2020UniversalDV). These computational tools provide a structured way to test hypotheses, refine mathematical models, and validate findings against real-world linguistic data (Hammarstrm2019OnCH).
While initially a topic of theoretical interest, the computational study of semantic change has intrinsic value for understanding language evolution and supporting linguistic research. Beyond its linguistic importance, it also informs certain computer science tasks that involve historical language or evolving terminology. For example, historical knowledge graphs such as Wikidata need to account for changes in entity and relation meanings over time to maintain consistency (Polleres2023HowDK). In digital humanities, tracking the semantic shift of key concepts (e.g., liberal, gay, computer) over centuries enables more accurate interpretation of historical texts. Similarly, domain-specific information retrieval can benefit from awareness of semantic drift — for example, the term cloud in technical literature has shifted significantly over the past decades.
Developing cost-effective tools that can handle linguistic variation and precisely characterize semantic changes is crucial for many applications. While the phenomena causing semantic change have been studied extensively in Linguistics (traugott2017semantic), providing deep analyses and typologies, and considerable effort has been devoted to detecting such changes by the computational linguistics community (Tahmasebi2018SurveyOC), to the best of our knowledge, few efforts have focused on formalizing the problem to characterize the type of change a word has undergone (Hengchen2021ChallengesFC).
@@
This survey addresses the need to understand and formalize the characterization of semantic change. Characterizing such changes will open new perspectives, particularly since we often lack the resources to study them in detail [^1] While NLP resources have hundreds of thousands of examples, datasets for semantic change generally cover fewer than a hundred words Kutuzov2022ContextualizedEF.. This is especially problematic in contexts where multiple word meanings are relevant, potentially leading to political or social misinterpretations. This survey’s objectives include:
@@
In this survey, we select categories that are both addressable by computational methods in a comparative setting (Tang2013SemanticCC) and well-established in the linguistic literature (campbell2013historical, p. 221-246), i.e., changes that can be systematically measured and are backed by theoretical analysis (Hammarstrm2019OnCH). For instance, euphemistic shifts often depend on social norms. When the word ‘weed’ acquired the sense of ‘cannabis’, this shift was motivated by the intention to obscure the literal meaning, given that cannabis carried negative or prohibited connotations within that particular cultural and historical context. Capturing this type of change requires not only knowledge of the lexical relation between ‘grass’ and ‘cannabis’, but also an understanding of external social and cultural factors.
Another example concerns cases in which a word’s sense becomes more general or more specific. Such changes cannot always be detected through direct comparison of the two senses, as the more general sense may still appear appropriate in both contexts, thus obscuring the shift. We focus on three factors of semantic change: orientation (i.e., whether a word’s meaning becomes more pejorative or positive), relation (i.e., whether a word’s meaning changes through figurative usage), and dimension (i.e., whether a word’s meaning becomes broader or more specific).
These poles of change also encode fundamental aspects of meaning: denotative (dimension), connotative (orientation), and cognitive (relation). Some of these properties, as we detail later, are crucial for understanding the intent of a sentence, for writing tools, to assist with translation, or to interact with users (e.g., in a social robot). Finally, this will open new opportunities to improve the reasoning capabilities of NLP tools (Danilevsky2020ASO).
@@ 2 Lexical Semantic Change
For the characterization of change, we adopt the mainstream linguistic typology for semantic change (traugott2017semantic): (i) broadening and narrowing, (ii) amelioration and pejoration, and (iii) metonymization and metaphorization. Before studying change characterization, we first review work on change identification. This is a well-developed topic that focuses on evaluating whether a word’s meaning has changed. Papers addressing change identification often state the need for further characterization of the type of change, but the state of computational change characterization remains unclear. This conclusion motivated the work presented in this paper and differentiates our survey from existing ones on LSC.
@@ 2.1 Related Works on Identification of Lexical Semantic Changes
Existing surveys on semantic change regularly cover detection but seldom focus on characterization*, which is the topic of this paper. In the context of identification, Tang2018ASO proposes splitting LSC analysis into four steps: obtaining and preparing a diachronic corpus, identifying word senses, modeling semantic change, and presenting and validating hypotheses. This empirical, rather than historical, paradigm enables the study of semantic change through scientific and statistical methods, allowing findings to be reproduced and tested—a key element for computational research.
@@
A second survey on computational approaches to LSC (Tahmasebi2018SurveyOC) focuses on word-level differentiation, sense differentiation, and lexical replacement. While it includes some types of change categories (e.g., metaphor, metonymy), the main change categories discussed are broadening, narrowing, and correlated types (split, join, birth).
@@
However, one recommendation from the authors is to conduct deeper studies on methods for automatically detecting other types of change and using them to build word histories. The authors provide an in-depth analysis of the linguistic literature grounding each domain, followed by evaluation strategies and applications. The difficulty of comparing approaches is highlighted and attributed to the absence of standard metrics and reference datasets.
@@
Kutuzov2018DiachronicWE survey word representation models and data for semantic change identification and describe a framework for comparing these methodologies. The authors focus on defining consistent terminology and coherent evaluation practices. In this work, they claim there’s still a gap in studying sub-classifications of change:
@@
“However, more detailed analysis of the nature of the shift is needed. […] This problem was to some degree addressed by Mitra2015AnAA, but much more work is certainly required to empirically test classification schemes proposed in much of the theoretical work […]”
@@
For the study of LSC and NLP in general, there are four main ways to represent word meaning: (i) frequency of co-occurrence, (ii) graphs of associated words or meanings, (iii) topic modeling for a list of meaningful terms, and (iv) word embeddings. Initial studies in the field employed graph- or frequency-based methods. Later approaches benefited from advances in the state of the art, replacing discrete methods with those based on static and contextual embeddings. Although contextual embeddings represented a massive improvement for a series of NLP tasks (Devlin2019BERTPO), the improvement for LSC did not follow this trend (Schlechtweg2020Semeval), showing that there is still room for improvement in this domain (Zhou2021FrequencybasedDI; Kutuzov2022ContextualizedEF).
@@
Other surveys in LSC investigate the phenomenon of change at the ontological level, trying to discover which laws govern the semantic evolution of words. For example, which kinds of words are more prone to broadening, given their current properties? In this context, Dubossarsky2017OuttaCL investigate existing evidence for claims such as (i) word frequency affects meaning change, (ii) a negative correlation exists between meaning change and prototypicality, and (iii) a positive correlation exists between meaning change and polysemy. The authors observed that laws proposed in previous studies are affected by model or data bias. When other investigators reduced the bias, the ‘laws’ either did not hold or were less precise than initially reported.
@@
Research in this field also delves into methodological issues, including the challenges of characterizing semantic shifts. As pointed out by Hengchen2021ChallengesFC, few studies detail the nature of these changes, mainly because it is difficult to model meaning from textual data. The paper further concludes that no particular measures for quantifying category change exist and that a consensus on a taxonomy for classifying types of change is required.
@@
In the work of Montanelli2023ASO, the authors review methodologies based on contextualized word embeddings. They categorize methodologies for training contextualized embeddings based on how they represent meaning, how time information is presented to the model, and what kind of learning procedure is performed. In this paper, the authors also discuss the importance of characterization:
@@
For this survey, we argue that semantic change characterization needs appropriate methods and tools to analyze each fine-grained category. We also defend that observing variations in a single property cannot describe all forms of LSC well. For instance, changes in the angle between word vectors (Dubossarsky2017OuttaCL) is not the best solution to identify the narrowing or broadening of a sense (Tahmasebi2018SurveyOC). Specific methods for identifying different types of change could lead to a better understanding of LSC than straightforward dissimilarity measures.
@@ 2.2 Objectives
Different terms refer to Lexical Semantic Change (LSC), such as semantic drift or semantic shift. In this paper, we distinguish between the identification of a change and the characterization of that change. The former refers to detecting that a word’s meaning has changed, while the latter refers to describing how it has changed. We formally differentiate these problems in Section 5.1. The analysis of work on the identification problem has been well covered by Tahmasebi2018SurveyOC, Kutuzov2018DiachronicWE, and Tang2018ASO.
Semantic change occurs from the need to convey more complex concepts and ideas with a restricted set of known words (HockJoseph2019Semantic). While some authors categorize changes by their cause (e.g., social, cultural), others use linguistic mechanisms (e.g., metaphor, metonymy) as the basis for classification. The focus of our literature review is on the characterization of change, where we look for common properties that allow for grouping identified semantic changes into categories. From a computational perspective, determining the cause of a semantic change is very complex to implement and, to the best of our knowledge, no computational approach has been implemented yet.
Frameworks have been proposed on how to track LSC, such as the Cambridge and McTaggart views (Tang2013SemanticCC), which did not initially consider computational methods. These approaches differ in how they use temporal information to identify changes: while the Cambridge view involves comparing two static corpora from different time periods, the McTaggart view involves tracking the evolution of a word’s usage over time. The Cambridge approach best suits the computational perspective, as the McTaggart approach would require historiographical work to acquire socio-cultural knowledge that goes beyond contextual analysis of word occurrences (Tang2013SemanticCC, p. 4-7).
@@
For this review, we assume the Cambridge setting for change analysis, using linguistic figures as a basis for change categories, which aligns with current state-of-the-art methods for computational LSC (Pivovarova2021RuShiftEvalAS). Given these considerations, we adopt the typology from traugott2017semantic, which is also predominant in the literature, as seen in koch2016meaning, HockJoseph2019Semantic, and campbell2013historical:
@@
Since the types in this typology form natural pairs that often involve similar computational challenges (e.g., broadening/narrowing requires sense counting; amelioration/pejoration requires sentiment analysis; metaphorization/metonymization requires a relatedness measure (koch2016meaning; Blank2003Polysemy; HockJoseph2019Semantic)), we group them into three overarching poles to better compare and categorize research. We only include approaches that compare corpora, as this typology is designed for diachronic studies rather than synchronic variation (e.g., across domains). Figure 1 illustrates the hierarchical taxonomy for LSC.
Figure 1: Taxonomy for the poles of Lexical Semantic Change, based on the work of koch2016meaning; Blank2003Polysemy; HockJoseph2019Semantic.
We assign studies that track sense broadening or narrowing to the dimension ($\mathcal{D}$) pole, where words gaining or losing senses see a change in their dimension value (see Section 5.2). This pole is self-complementary, as the change in the dimension’s value addresses both broadening and narrowing. A change in dimension could signal, for example, that a word requires more careful translation (Mohammed2009PolysemyAA).
We group changes related to metaphorization or metonymization into the relation ($\mathcal{R}$) pole. These changes occur when a new sense develops from an existing one through a figurative link (see Section 5.3). Relations like hypernymy, hyponymy, and synonymy are also factors for semantic change (koch2016meaning; HockJoseph2019Semantic) as they involve a conceptual relation and can potentially be encompassed in this pole.
For the orientation ($\mathcal{O}$) pole, we group studies that identify amelioration, pejoration, or any perceptual change in the sentiment associated with a word (e.g., criticism, compliment). This pole also encompasses many forms of change in perception, as some studies suggest that concepts like ‘taboo’ also affect changes in word usage (koch2016meaning).
@@
Figure 2: Change in meaning and orientation for the word awful. In the left side, we reproduce the figure from Hamilton2016DiachronicWE that shows the evolution of the word ‘awful’ in the embedding space. In the right side, we present the hypothetical function ($\mathfrak{f}$) for this word over time.

### v5

# Survey in Characterizing Semantic Change

@@ Abstract
Living languages continuously evolve to reflect the cultural changes of human societies. This evolution manifests through neologisms (new words) or the semantic change of existing words (new meanings for existing words). Understanding the meaning of words is vital for interpreting texts from different cultures (regionalisms or slang), domains (e.g., technical terms), or time periods. In computer science, this phenomenon is relevant to computational linguistics tasks such as machine translation, information retrieval, and question answering. Semantic change can impact the performance of these applications, making it important to understand and characterize these changes formally. This problem has recently attracted significant attention from the computational linguistics community. Several approaches can detect semantic changes with good precision, but more effort is needed to characterize how word meanings change and to determine how to mitigate the impact of this phenomenon. This survey provides a comprehensive overview of existing approaches to the characterization of semantic change. We also formally define three classes of characterization: change in dimension (whether a word’s meaning becomes broader or narrower), change in orientation (whether a word acquires a more pejorative or ameliorative sense), and change in relation (whether a word is used in a new figurative context, such as a metaphor or metonymy). We demonstrate the applicability of this formalism on existing corpora, summarize the key aspects of selected publications, and discuss current needs and trends in research on semantic change characterization.
@@ 1 Introduction
Language is one of the most dynamic skills humans have acquired, constantly evolving to adapt communication to new purposes and needs (Pinker, 2003, p. 13-23). While no single unified theory explains its origins (Allan, 2013, p. 13-52), language has clearly undergone significant modifications throughout human history. One of the prevailing theories of language evolution is based on the assumption of “continuity of change” (Rilling et al., 2012).
This theory posits that language, being too complex to have appeared suddenly, must have evolved gradually from earlier pre-linguistic systems used by our ancestors (Pinker and Bloom, 1990, p. 725-726). In this context, comprehending language change contributes to the fields of history, sociology, linguistics, philosophy, and Natural Language Processing (NLP) (Campbell, 2013, p. 1-8), as it is essential to understanding the origins of language, cultural development, grammatical theories, and the relationship between language and thought (Sherwood et al., 2008, p. 426–454).
@@
  • Phonological Changes: Changes in the pronunciation and sound systems of a language. For instance, the pronunciation of vowel sounds changed significantly due to the Great Vowel Shift in Medieval English (Nevalainen and Traugott, 2016, p. 367).
@@
  • Syntactic Changes: Changes in grammar and word order. For example, Modern English emphasizes word order and auxiliary verbs more than the intricate inflectional system of Old English (Nevalainen and Traugott, 2016, p. 616).
@@
While phonological and morphological changes are crucial, semantics plays a pivotal role in language evolution. Semantic change reflects how words adapt to evolving cultural, technological, and societal contexts, creating new meanings and preserving language’s fundamental purpose of effective communication. Despite being a recurring topic in scientific research, semantic change remains less understood than other forms of linguistic evolution due to its often subtle nature compared to more observable shifts, like those resulting from grammar reforms (Allan and Robinson, 2012, p. 2-3).
@@
The advent of computational linguistics and advanced models marked a major shift in the field. With vast digital corpora and efficient algorithms, researchers began to apply statistical methods from natural language processing to analyze text automatically. This marked a transition from purely observational methods to data-driven statistical analysis (Allan and Robinson, 2012, p. 161-164).
With access to large text corpora and powerful algorithms, linguists can now detect subtle patterns in language with greater accuracy and efficiency. This allows for the study of languages with limited prior linguistic knowledge (Conneau et al., 2017; Wu and Dredze, 2019), given sufficient data, and enables comparison of multiple languages to test for universal properties in latent space (Vulic et al., 2020; Wang et al., 2022).
By adopting a model-based approach, linguists can now explore complex language phenomena across different cultures and backgrounds (Nivre et al., 2020). These computational tools provide a structured way to test hypotheses, refine mathematical models, and validate findings against real-world linguistic data (Hammarström et al., 2019).
While initially a topic of theoretical interest, the computational study of semantic change has intrinsic value for understanding language evolution and supporting linguistic research. Beyond its linguistic importance, it also informs certain computer science tasks that involve historical language or evolving terminology. For example, historical knowledge graphs such as Wikidata need to account for changes in entity and relation meanings over time to maintain consistency (Polleres et al., 2023). In digital humanities, tracking the semantic shift of key concepts (e.g., liberal, gay, computer) over centuries enables more accurate interpretation of historical texts. Similarly, domain-specific information retrieval can benefit from awareness of semantic drift — for example, the term cloud in technical literature has shifted significantly over the past decades.
Developing cost-effective tools that can handle linguistic variation and precisely characterize semantic changes is crucial for many applications. While the phenomena causing semantic change have been studied extensively in Linguistics (Traugott, 2017), providing deep analyses and typologies, and considerable effort has been devoted to detecting such changes by the computational linguistics community (Tahmasebi et al., 2018), to the best of our knowledge, few efforts have focused on formalizing the problem to characterize the type of change a word has undergone (Hengchen et al., 2021).
@@
This survey addresses the need to understand and formalize the characterization of semantic change. Characterizing such changes will open new perspectives, particularly since we often lack the resources to study them in detail [^1] While NLP resources have hundreds of thousands of examples, datasets for semantic change generally cover fewer than a hundred words Kutuzov et al. (2022).. This is especially problematic in contexts where multiple word meanings are relevant, potentially leading to political or social misinterpretations. This survey’s objectives include:
@@
In this survey, we select categories that are both addressable by computational methods in a comparative setting (Tang et al., 2013) and well-established in the linguistic literature (Campbell, 2013, p. 221-246), i.e., changes that can be systematically measured and are backed by theoretical analysis (Hammarström et al., 2019). For instance, euphemistic shifts often depend on social norms. When the word ‘weed’ acquired the sense of ‘cannabis’, this shift was motivated by the intention to obscure the literal meaning, given that cannabis carried negative or prohibited connotations within that particular cultural and historical context. Capturing this type of change requires not only knowledge of the lexical relation between ‘grass’ and ‘cannabis’, but also an understanding of external social and cultural factors.
Another example concerns cases in which a word’s sense becomes more general or more specific. Such changes cannot always be detected through direct comparison of the two senses, as the more general sense may still appear appropriate in both contexts, thus obscuring the shift. We focus on three factors of semantic change: orientation (i.e., whether a word’s meaning becomes more pejorative or positive, for example, the word awful acquired a negative meaning), relation (i.e., whether a word’s meaning changes through figurative usage, ‘head’ used to address the director of a company), and dimension (i.e., whether a word’s meaning becomes broader or more specific, ‘cell’ used to mean only a room in prison).
These poles of change also encode fundamental aspects of meaning: denotative (dimension), connotative (orientation), and cognitive (relation). Some of these properties, as we detail later, are crucial for understanding the intent of a sentence, for writing tools, to assist with translation, or to interact with users (e.g., in a social robot). Finally, this will open new opportunities to improve the reasoning capabilities of NLP tools (Danilevsky et al., 2020).
@@ 2 Lexical Semantic Change
For the characterization of change, we adopt the mainstream linguistic typology for semantic change (Traugott, 2017): (i) broadening and narrowing, (ii) amelioration and pejoration, and (iii) metonymization and metaphorization. Before studying change characterization, we first review work on change identification. This is a well-developed topic that focuses on evaluating whether a word’s meaning has changed. Papers addressing change identification often state the need for further characterization of the type of change, but the state of computational change characterization remains unclear. This conclusion motivated the work presented in this paper and differentiates our survey from existing ones on LSC.
@@ 2.1 Related Works on Identification of Lexical Semantic Changes
Existing surveys on semantic change regularly cover detection but seldom focus on characterization, which is the topic of this paper. In the context of identification, Tang (2018) proposes splitting LSC analysis into four steps: obtaining and preparing a diachronic corpus, identifying word senses, modeling semantic change, and presenting and validating hypotheses. This empirical, rather than historical, paradigm enables the study of semantic change through scientific and statistical methods, allowing findings to be reproduced and tested—a key element for computational research.
@@
A second survey on computational approaches to LSC (Tahmasebi et al., 2018) focuses on word-level differentiation, sense differentiation, and lexical replacement. While it includes some types of change categories (e.g., metaphor, metonymy), the main change categories discussed are broadening, narrowing, and correlated types (split, join, birth).
@@
However, one recommendation from the authors is to conduct deeper studies on methods for automatically detecting other types of change and using them to build word histories. The authors provide an in-depth analysis of the linguistic literature grounding each domain, followed by evaluation strategies and applications. The difficulty of comparing approaches is highlighted and attributed to the absence of standard metrics and reference datasets, but also to the inherent challenges of the task itself.
@@
Kutuzov et al. (2018) survey word representation models and data for semantic change identification and describe a framework for comparing these methodologies. The authors focus on defining consistent terminology and coherent evaluation practices. In this work, they claim there’s still a gap in studying sub-classifications of change:
@@
“However, more detailed analysis of the nature of the shift is needed. […] This problem was to some degree addressed by Mitra et al. (2015), but much more work is certainly required to empirically test classification schemes proposed in much of the theoretical work […]”
@@
For the study of LSC and NLP in general, there are four main ways to represent word meaning: (i) frequency of co-occurrence, (ii) graphs of associated words or meanings, (iii) topic modeling for a list of meaningful terms, and (iv) word embeddings. Initial studies in the field employed graph- or frequency-based methods. Later approaches benefited from advances in the state of the art, replacing discrete methods with those based on static and contextual embeddings. Although contextual embeddings represented a massive improvement for a series of NLP tasks (Devlin et al., 2019), the improvement for LSC did not follow this trend (Schlechtweg et al., 2020), showing that there is still room for improvement in this domain (Zhou et al., 2021; Kutuzov et al., 2022).
@@
Other surveys in LSC investigate the phenomenon of change at the ontological level, trying to discover which laws govern the semantic evolution of words. For example, which kinds of words are more prone to broadening, given their current properties? In this context, Dubossarsky et al. (2017) investigate existing evidence for claims such as (i) word frequency affects meaning change, (ii) a negative correlation exists between meaning change and prototypicality, and (iii) a positive correlation exists between meaning change and polysemy. The authors observed that laws proposed in previous studies are affected by model or data bias. When other investigators reduced the bias, the ‘laws’ either did not hold or were less precise than initially reported.
@@
Research in this field also delves into methodological issues, including the challenges of characterizing semantic shifts. As pointed out by Hengchen et al. (2021), few studies detail the nature of these changes, mainly because it is difficult to model meaning from textual data. The paper further concludes that no particular measures for quantifying category change exist and that a consensus on a taxonomy for classifying types of change is required.
@@
In the work of Montanelli and Periti (2023), the authors review methodologies based on contextualized word embeddings. They categorize methodologies for training contextualized embeddings based on how they represent meaning, how time information is presented to the model, and what kind of learning procedure is performed. In this paper, the authors also discuss the importance of characterization:
@@
For this survey, we argue that semantic change characterization needs appropriate methods and tools to analyze each fine-grained category. We also defend that observing variations in a single property cannot describe all forms of LSC well. For instance, changes in the angle between word vectors (Dubossarsky et al., 2017) is not the best solution to identify the narrowing or broadening of a sense (Tahmasebi et al., 2018). Specific methods for identifying different types of change could lead to a better understanding of LSC than straightforward dissimilarity measures.
@@ 2.2 Objectives
Different terms refer to Lexical Semantic Change (LSC), such as semantic drift or semantic shift. In this paper, we distinguish between the identification of a change and the characterization of that change. The former refers to detecting that a word’s meaning has changed, while the latter refers to describing how it has changed. We formally differentiate these problems in Section 5.1. The analysis of work on the identification problem has been well covered by Tahmasebi et al. (2018), Kutuzov et al. (2018), and Tang (2018).
Semantic change occurs from the need to convey more complex concepts and ideas with a restricted set of known words (Hock and Joseph, 2019). While some authors categorize changes by their cause (e.g., social, cultural), others use linguistic mechanisms (e.g., metaphor, metonymy) as the basis for classification. The focus of our literature review is on the characterization of change, where we look for common properties that allow for grouping identified semantic changes into categories. From a computational perspective, determining the cause of a semantic change is very complex to implement and, to the best of our knowledge, no computational approach has been implemented yet.
Frameworks have been proposed on how to track LSC, such as the Cambridge and McTaggart views (Tang et al., 2013), which did not initially consider computational methods. These approaches differ in how they use temporal information to identify changes: while the Cambridge view involves comparing two static corpora from different time periods, the McTaggart view involves tracking the evolution of a word’s usage over time. The Cambridge approach best suits the computational perspective, as the McTaggart approach would require historiographical work to acquire socio-cultural knowledge that goes beyond contextual analysis of word occurrences (Tang et al., 2013, p. 4-7).
@@
For this review, we assume the Cambridge setting for change analysis, using linguistic figures as a basis for change categories, which aligns with current state-of-the-art methods for computational LSC (Pivovarova and Kutuzov, 2021). Given these considerations, we adopt the typology from Traugott (2017), which is also predominant in the literature, as seen in Koch (2016, p. 21-66), Hock and Joseph (2019, p. 189-222), and Campbell (2013, p. 221-246):
@@
Since the types in this typology form natural pairs that often involve similar computational challenges (e.g., broadening/narrowing requires sense counting; amelioration/pejoration requires sentiment analysis; metaphorization/metonymization requires a relatedness measure (Koch, 2016; Blank and others, 2003; Hock and Joseph, 2019)), we group them into three overarching poles to better compare and categorize research. We only include approaches that compare corpora, as this typology is designed for diachronic studies rather than synchronic variation (e.g., across domains). Figure 1 illustrates the hierarchical taxonomy for LSC.
Figure 1: Taxonomy for the poles of Lexical Semantic Change, based on the work of Koch (2016); Blank and others (2003); Hock and Joseph (2019).
We assign studies that track sense broadening or narrowing to the dimension ($\mathcal{D}$) pole, where words gaining or losing senses see a change in their dimension value (see Section 5.2). This pole is self-complementary, as the change in the dimension’s value addresses both broadening and narrowing. A change in dimension could signal, for example, that a word requires more careful translation (Mohammed, 2009).
We group changes related to metaphorization or metonymization into the relation ($\mathcal{R}$) pole. These changes occur when a new sense develops from an existing one through a figurative link (see Section 5.3). Relations like hypernymy, hyponymy, and synonymy are also factors for semantic change (Koch, 2016; Hock and Joseph, 2019) as they involve a conceptual relation and can potentially be encompassed in this pole.
For the orientation ($\mathcal{O}$) pole, we group studies that identify amelioration, pejoration, or any perceptual change in the sentiment associated with a word (e.g., criticism, compliment). This pole also encompasses many forms of change in perception, as some studies suggest that concepts like ‘taboo’ also affect changes in word usage (Koch, 2016).
@@
Figure 2: Change in meaning and orientation for the word awful. In the left side, we reproduce the figure from Hamilton et al. (2016b) that shows the evolution of the word ‘awful’ in the embedding space. In the right side, we present the hypothetical function ($\mathfrak{f}$) for this word over time.